Frank invest a certain amount of money at 5% intrest and $1360 more than that amount at 7%. His total yearly interest was $241.60. How much did she invest at each rate?

Answer by addingup(3677)

Let the amount at 5% be x. Then, the amount at 7% will be x+1360.
0.05x+0.07(x+1360) = 241.60
0.05x+0.07x+95.20 = 241.60
0.12x = 146.40
x = 1220 is invested at 5% and
1220+1360 = 2580 is invested at 7%
:
John